2 Question 2 Tries remaining: 2and the price of a unit of paper is $109. Calculate the profit-maximizing quantity of paper. Points out of 6.25 (Round to the nearest two decimal places if necessary.) Suppose that paper factory had a total cost function given by: TC 588 +q+3q. The paper factory is a price taker Flag question Answer: Check
Answer #1

TC = 588 + q + 3q2

Marginal cost (MC) = d/dQ (TC) = d/dQ (588 + q + 3q2) = 6q

For a price taking firm, price is marginal revenue (MR). So, MR = $109

At profit maximizing quantity MR = MC

i.e. 109 = 6q

or, q = 109/6 = 18.17

Answer: 18.17